OpenStudy (anonymous):

Steve travels 4 times as fast as Fred. Traveling in opposite directions, they are 105 miles apart at 3 hours find the rates of travel I know rt=d

OpenStudy (anonymous):

After 3 hours that is

OpenStudy (anonymous):

The combined rate is 5r because 4r - (-r) = 5r. 5r = d/3. You have d, so just solve for r.

OpenStudy (anonymous):

Fred's rate will be "r" and Steve's rate will be "4r"

OpenStudy (anonymous):

Ok so how do I set it up to solve

OpenStudy (anonymous):

?..

OpenStudy (anonymous):

I already set up the equation for you. 5r = d/3. First substitute in your "d" value.

OpenStudy (anonymous):

5r=3 ?? What

OpenStudy (anonymous):

Not 5r equal 3. I wrote that 5r = distance divided by 3. d is distance and it is part of the equation.

OpenStudy (anonymous):

Ok sorry that took me a while to understand lol got it thanks

OpenStudy (anonymous):

You're welcome! And medals are nice!